# Simulated host galaxy analogs of high-z quasars observed with JWST

### v1

@@ Abstract
The deconvolving of the quasar from its host galaxy’s stellar light requires careful fitting of the point source and galactic light profile. This ensures an accurate estimation of parameters such as the magnitude, mass, and radius. Ding et al. (2023) used the galight software (Ding et al., 2020) to decouple the stellar light from the quasar by modeling the quasar as a two dimensional PSF and Sérsic profile of the host galaxy. This is done separately in both the F356W and F150W filters. This was successful except for J2255+0251 which was only detected in F356W.
@@
Stone et al. (2023b) also detected the host of J2239+0207 at by scaling the PSF to match the quasar flux and finding a discrepancy in the difference between the two. We do not include this detection’s measurements in our paper to stay consistent with detections that are performed with Bayesian point source removal software, i.e., psfmc (Mechtley, 2014) and galight software (Ding et al., 2020). With non-Bayesian methods, Stone et al. (2023a) set upper limits on the galaxy flux from five high-z quasars that were not detected with JWST. These quasar host detections and non-detections have provided a first glimpse into measuring the black hole to stellar mass relationship in high-z quasars. The high-z Magorrian relation has since been seen to show a higher scatter than at low-z (Stone et al., 2023a). For example, the EIGER quasars are reported to have black hole to stellar mass ratios that are up to 2 dex higher than the local relation (Yue et al., 2023).
@@
In this paper, we examine the measured properties of J2255+0251, J2236+0032, and the EIGER quasars with detected hosts within the distribution of galaxies with black holes simulated in the BlueTides simulation. In section 2, we provide a brief description of the BlueTides simulation. In section 3, we describe the specifics of converting the observed JWST parameters to comparable quantities in BlueTides. We then quantify our comparison between these quasars and the predictions of the BlueTides simulation. In section 4, we describe our methods for generating mock images and spectra of the BlueTides analogs of J2255+0251 and J2236+0032 and also performing a point source removal. We compare the results from the mock spectra, images, and point source removal in section 5. We make predictions for quasar host detectability in upcoming observations of the quasar analogs with the followup filters and exposure times in section 6. Lastly in section 7, we summarize our conclusions and provide future outlooks on upcoming JWST quasar measurements. We assume a cosmology with parameters from WMAP9 (Hinshaw et al. 2013) to maintain consistency with the BlueTides simulation as follows, , , , , , and . All magnitudes are expressed in AB magnitude.
@@
Parameters with an were calculated from observed parameters using methods described in section 3.1.
@@
Parameters with an were calculated from observed parameters using methods described in section 3.1.
@@ 3.3 Stellar mass and SFR
Figures 1(b) through 2(a) show the joint distribution of black hole masses and luminosity with stellar mass. The stellar mass was measured as and for J2255+0251 and J2236+0032, respectively (Ding et al., 2023). The black hole masses and stellar masses of J2255+0251 and J2236+0032 follow the black hole mass to stellar mass ratio. However, the EIGER quasars lie 1-2 dex above the Kormendy & Ho (2013) relation and do no lie within the BlueTides distribution. The EIGER quasars have bolometric luminosities that are brighter than the BlueTides quasars which may bias the comparison. The uncertainty and difficulty of removing the quasar from its host (as described in section 4.3) could lead to biased stellar mass measurements. Understanding this observational bias will be the goal of future work.
In Figure 2(b), we show where the quasars would lie on the star formation rate versus bolometric luminosity distribution, allowing us to estimate their SFR. We anticipate a star formation rate between approximate for both quasars. For the more luminous EIGER quasars, we can place an approximate lower limit of . Followup measurements in the sub-mm with instruments such as ALMA of the H emission line with JWST will directly measure the SFR. This shows the potential of using high-z simulations to both constrain and support future quasar observations.
@@ 4.1 BlueTides analog selection
We adjust the range to explore for each property to ensure that we find at least one analog per observed quasar. Our analogs are within of the measured bolometric luminosities for both quasars. They are within and of the measured black hole masses and and of the measured stellar masses for J2236+0032 and J2255+0251, respectively. The J2255+0251 BlueTides analog is within the error bars of the measured galaxy radius, but the J2236+0032 analog is about larger in galaxy radius than the observed host. In Figure 1(a), the J2255+0251 analog is seen with a black hole mass and bolometric luminosity nearly matching the observed quasar host. For J2255+0251, we choose the BlueTides galaxy with the closest properties as our analog, but for J2236+0032, we select the brightest quasar from approximately 5 analogs to remain closer to the magnitude of the detected host in Ding et al. (2023). This allows for comparison of the point source subtraction between two analogs with varying magnitudes for J2236+0032. With a BlueTides galaxy analog for both J2236+0032 and J2255+0251, we are able to make comparisons and predictions between BlueTides and these high-z quasars.
@@ 4.4 Point source removal
After the creation of mock images, we recover the properties of the host galaxy and quasar from the imaged galaxies using the software psfmc to do point source extraction. psfmc was written specifically for quasar subtraction unlike other algorithms that perform similar 2D surface brightness fitting such as galfit in Peng et al. (2002). psfmc uses the Affine Invariant Markov chain Monte Carlo (MCMC) ensemble sampler implemented in emcee (Foreman-Mackey et al., 2013). The posterior sampled includes priors specified by the user and the following Gaussian joint likelihood for all pixels
@@
The output of psfmc contains the posteriors of all parameters and the final galaxy image with the quasar subtracted. The intermediate images derived are the raw model and the convolved model. The raw model contains the highest probability galaxy and quasar surface brightness distribution. We convolve the raw model with the PSF to form the convolved model which is subtracted from the input observed image to determine the residuals. The third column of the panels labeled "model" in figures 6(a) through 6(d) show the convolved model.
@@
We initiate 64 walkers and discard the initial 200-400 samples to allow proper convergence of our sampler. The values of our results are drawn from the last 1k sample ensuring accurate parameter error estimation. The mock images are shown and compared qualitatively in section 5.
5 Comparing BlueTides Mock Observations to the SHELLQs Quasars
5.1 Mock Spectra of BlueTides Analogs
We make redshifted mock spectra for the JWST analogs of J2255+0251 and J2236+0032. We infer the SED in the measured two-band photometry as described in section 4.3. The magnitudes and flux ratios for both the observed and simulated quasars are summarized in Table 3. Notably, the BlueTides host analogs have magnitudes broadly similar to the magnitudes of the observed quasar hosts. The upper limit for the host of J2255+0251 in F150W falls significantly below the magnitude of the BlueTides analog. For the filters where hosts were detected, this provides evidence that the analogs chosen are representative of the observed quasar hosts.
In Figure 5, we plot the galaxy spectra for both analogs to compare to Extended Data Figure 5 in Ding et al. (2023). Although we do not perform a quantitative comparison, the spectra look significantly different on each side of the Balmer break. Blueward of the Balmer break, the analogs show a much steeper spectra than the Ding et al. (2023) median template from SED inference, instead being more consistent with their template corresponding to the 1 upper limits. This results in the significantly higher flux in both BlueTides analogs in the F150W filter. This different spectral shape likely implies a different age, metallicity, and/or dust attenuation for our galaxy than the median estimated by Ding et al. (2023).
At wavelengths longer than the Balmer break, there is significantly less reddening, which could be due to the different dust model we assume. For both observed quasar hosts, the spectra fit must be lower in F150W and higher in F356W to match the observed magnitudes. This may indicate that the observed host brightness was underestimated, due to the difficulty of the quasar subtraction. These mock spectra show the diversity of emission lines potentially present in the high-z quasar hosts which are much more difficult to detect in high-z quasar systems compared to the quasar spectra.
J2255 | F356W | 24.2 | 12.5 | ||
J2255 | F150W | 24.8 | 14.7 | ||
J2236 | F356W | 23.5 | 18.9 | ||
J2236 | F150W | 24.6 | 15.0 |
5.2 Mock Rest Frame Images of BlueTides Analogs
With the intrinsic properties simulated in BlueTides, we also create rest frame images of galaxies. These provide a glimpse into the evolution of detected hosts of the Ding et al. (2023) and other SHELLQs quasars that is unachievable with observations. We use SynthObs to generate rest frame images of the BlueTides analogs. These images do not have the observational effects of the other mock images, such as noise and PSF effects, and have a much higher resolution. We show the stellar mass distribution and recent star formation in Figure 6. We also show images in idealized top hat filters in the far UV, visible (V), and infrared (H) as well as the stacked RBG version. These images provide a higher resolution snapshot of the star particles in the BlueTides analog than the blurred redshifted images shown in Figures 6(a) through 7(b). The galaxies are notably compact in the left three panels in stellar mass, SFR, and intrinsic FUV. The intrinsic FUV panel traces the star formation and stellar mass. This is in line with results from Marshall et al. (2022a) which predicted that galaxies hosting quasars are more compact intrinsically relative to the entire BlueTides sample. When we include dust attenuation in the four right panels of Figure 6, we see a flatter emission profile as predicted in Marshall et al. (2022a). This results in the appearance of an extended galaxy. The ring around the center of these galaxies shows a region where the dust is particularly strong.
@@ 5.3 Mock Images of BlueTides Analogs
In Figures 6(a) through 7(b), we show six panels of our mock observed images, depicting I) the BlueTides galaxy without a quasar imaged, II) the same BlueTides galaxy with a quasar imaged, III) the convolved model from psfmc described in section 4.3, IV) the host galaxy image with the quasar subtracted via psfmc, V) a normalized residual between the data and point source subtracted image from psfmc (i.e., (data - model) / ), and VI) the detected host JWST image from Ding et al. (2023). With psfmc, we are successful in extracting the 2D galactic brightness distribution and quasar parameters for all of our imaged BlueTides galaxies. This is promising as the analogs are particularly faint in the F150W filter. The actual host (galaxy only image) versus quasar subtracted (data - quasar image) all recover the underlying host with moderate visible error. In Figures 6(b) and 6(c), the center of the host is missing flux. Each quasar subtracted host also features subtle to significant morphological changes from the actual host image. Through the normalized residuals in the fourth panel, we can quantitatively compare the data + quasar versus galaxy only images which ideally would be the same. There is a slight underextraction and overextraction around the core area containing the point source. This is an artifact of the point source removal and should be considered with point source subtraction on real images.
In Figures 6(a) through 6(c), the J2236+0032 and J2255+0251 host analogs are similar to the detected hosts in Ding et al. (2023) in all filters. In Figure 6(d), the J2255+0251 host analog detection in F150W is shown alongside the non-detection in Ding et al. (2023). Notably, this is also the dimmest host analog detection in our sample.
@@ 5.4 Measured Properties in Mock Images
Qualitatively, the imaged galaxies look similar to the detected galaxies in Ding et al. (2023), with similar shapes and compactness. However for most of these mock images, our point source subtraction algorithm described in section 4.3 creates moderate to large errors in predicted radii and magnitude when extracting a point source. We compare the Sérsic radii and magnitudes quantitatively for the imaged galaxy with and without the point source. We extract both Sérsic radii and magnitudes using the MCMC framework in psfmc to maintain consistency. The Sérsic profile is defined as
where is the half light radius we define as the Sérsic radius, n is the Sérsic index corresponding to increasing steepness in the log brightness and log radius space, , and is the intensity at the half light radius. We tabulate the mean and one sigma value from our sampled posteriors for each Sérsic radius in Table 4 for both the original galaxy before adding the point source and the radius estimated after point source subtraction. We find that the results do not always agree to one sigma for both the original and point source subtracted galaxies indicating that point source subtraction is introducing bias.
@@ 6 Detection predictions for J2255+0251 and J2236+0032
Follow up observations with JWST of J2255+0251 and J2236+0032 (detected in both F150W and F356W by JWST) will allow further constraints on the properties of their hosts. In JWST Cycle 2 GO 3859 (PI: Onoue), these quasars will be followed up in five NIRCam filters. In Table 5, the filters and exposure times for this proposal are shown. In Figure 10, we show the mock images of galaxies without an added quasar for these filters and exposure times. We also extend the exposure time of the original filters (F356W and F150W) to 10ks in Figure 11 for comparison to the 3.1ks exposure times shown in figures 6(a) through 6(d). We supersample in all filters with wavelengths larger than (filters above F200W) to maintain consistency with our mock imaging in section 4.

### v3

@@ Abstract
The deconvolving of the quasar from its host galaxy’s stellar light requires careful fitting of the point source and galactic light profile. This ensures an accurate estimation of parameters such as the Sérsic magnitude and radius. Ding et al. (2023) used the galight software (Ding et al., 2020) to decouple the stellar light from the quasar by modeling the quasar as a two dimensional PSF and Sérsic profile of the host galaxy. The Sérsic profile is defined as
where is the half light radius we define as the Sérsic radius, n is the Sérsic index corresponding to increasing steepness in the log brightness and log radius space, , and is the intensity at the half light radius. This is done separately in both the F356W and F150W filters. This was successful except for J2255+0251 which was only detected in F356W.
@@
Stone et al. (2023) also detected the host of J2239+0207 at by scaling the PSF to match the quasar flux and finding a discrepancy in the difference between the two. We do not include this detection’s measurements in our paper to stay consistent with detections that are performed with Bayesian point source removal software, i.e., psfmc (Mechtley, 2014) and galight software (Ding et al., 2020). With non-Bayesian methods, Stone et al. (2024) set upper limits on the galaxy flux from five high-z quasars that were not detected with JWST. These quasar host detections and non-detections have provided a first glimpse into measuring the black hole to stellar mass relationship in high-z quasars. Although Ding et al. (2023)’s black hole mass measurements seem to be in line with the low-z relation, many other observations have measured overmassive black holes (Yue et al., 2023; Stone et al., 2024). For example, the EIGER quasars are reported to have black hole to stellar mass ratios that are up to 2 dex higher than the local relation (Yue et al., 2023).
@@
In this paper, we examine the measured properties of J2255+0251, J2236+0032, and the EIGER quasars with detected hosts within the distribution of galaxies with black holes simulated in the BlueTides simulation. In section 2, we provide a brief description of the BlueTides simulation. In section 3, we describe the specifics of converting the observed JWST parameters to comparable quantities in BlueTides. We then quantify our comparison between these quasars and the predictions of the BlueTides simulation. In section 4, we describe our methods for generating mock images and spectra of the BlueTides analogs of J2255+0251 and J2236+0032 and also performing a point source removal. We compare the results from the mock spectra, images, and point source removal in section 5. We make predictions for quasar host detectability in upcoming observations of the quasar analogs with the followup filters and exposure times in section 6. Lastly in section 7, we summarize our conclusions and provide future outlooks on upcoming JWST quasar measurements. We assume a cosmology with parameters from WMAP9 (Hinshaw et al. 2013) to as follows, , , , , , and . We express all magnitudes in AB magnitude.
@@
Parameters with an ∗ were calculated from observed parameters using methods described in section 3.1.
@@
Parameters with an ∗ were calculated from observed parameters using methods described in section 3.1.
@@ 3.3 Stellar mass and SFR
Figures 2(b) through 3(a) show the joint distribution of black hole masses and luminosity with stellar mass. The stellar mass was measured as and for J2255+0251 and J2236+0032, respectively (Ding et al., 2023). The black hole masses and stellar masses of J2255+0251 and J2236+0032 follow the black hole mass to stellar mass ratio. However, the EIGER quasars lie 1-2 dex above the Kormendy & Ho (2013) relation and do no lie within the BlueTides distribution. The EIGER quasars have bolometric luminosities that are brighter than the BlueTides quasars which may bias the comparison. The uncertainty and difficulty of removing the quasar from its host (as described in section 4.3) could lead to biased stellar mass measurements. Understanding this observational bias will be the goal of future work.
In Figure 3(b), we show where the quasars would lie on the star formation rate versus bolometric luminosity distribution. This allows us to estimate their SFR using the BlueTides distribution of quasars. We anticipate a star formation rate between approximate for both quasars. For the more luminous EIGER quasars, we can place an approximate lower limit of . Followup measurements in the sub-mm with instruments such as ALMA of the H emission line with JWST will directly measure the SFR. This shows the potential of using high-z simulations to both constrain and support future quasar observations.
@@ 4.1 BlueTides analog selection
We adjust the range to explore for each property to ensure that we find at least one analog per observed quasar. Our analogs are within 0.4 dex of the measured bolometric luminosities for both quasars. They are within 0.4 and 0.01 dex of the measured black hole masses and 0.2 and 0.01 dex of the measured stellar masses for J2236+0032 and J2255+0251, respectively. The J2255+0251 BlueTides analog is within the error bars of the measured galaxy radius, but the J2236+0032 analog is about 0.7 kpc larger in galaxy radius than the observed host. In Figure 2(a), the J2255+0251 analog is seen with a black hole mass and bolometric luminosity nearly matching the observed quasar host. For J2255+0251, we choose the BlueTides galaxy with the closest properties as our analog, but for J2236+0032, we select the brightest quasar from approximately 5 analogs to remain closer to the magnitude of the detected host in Ding et al. (2023). This allows for comparison of the point source subtraction between two analogs with varying magnitudes for J2236+0032. With a BlueTides galaxy analog for both J2236+0032 and J2255+0251, we are able to make comparisons and predictions between BlueTides and these high-z quasars.
@@ 4.3 Mock images and spectroscopy
5 Comparing BlueTides Mock Observations to the SHELLQs Quasars
5.1 Mock Spectra of BlueTides Analogs
We make redshifted mock spectra for the JWST analogs of J2255+0251 and J2236+0032 and 15 other close analogs with similar bolometric luminosities (some of their mock images can be seen in Figure 9). We infer the SED in the measured two-band photometry as described in section 4.3. The magnitudes and flux ratios for both the observed and simulated quasars are summarized in Table 3. Notably, the BlueTides host analogs have magnitudes broadly similar to the magnitudes of the observed quasar hosts. The upper limit for the host of J2255+0251 in F150W falls significantly below the magnitude of the BlueTides analog. For the filters where hosts were detected, this provides evidence that the primary analogs chosen are representative of the observed quasar hosts. For both observed quasar hosts, the spectra fit must be lower in F150W and higher in F356W to match the observed magnitudes. This may indicate that the observed host brightness was underestimated, due to the difficulty of the quasar subtraction.
In Figure 5, we plot the galaxy spectra for both the primary analogs and additional analogs to compare to Extended Data Figure 5 in Ding et al. (2023). The mock spectra also look similar to high-z galaxy spectra observed with JWST in Maiolino et al. (2023) and Boyett et al. (2024). This indicates reasonably accurate spectral modeling with SynthObs of the analogs.
Each analog has a similar spectral shape. The increasing amplitude of the spectra is roughly correlated with stellar mass but also depends on differences in average age, metallicity, and metal surface density for stars in the galaxies. All analog mock spectra also show the diversity of emission lines potentially present in the high-z quasar hosts which are much more difficult to detect in high-z quasar systems compared to the quasar spectra.
Although we do not perform a quantitative comparison, the spectra of the galaxy analogs look significantly different than the median template from SED inference in Ding et al. (2023). Blueward of the Balmer break, the analogs show a much steeper spectra, instead being more consistent with their template corresponding to the 1 upper limits. This results in the significantly higher flux in both primary BlueTides analogs in the F150W filter than the observed hosts. We predict this difference in spectral shape between all analogs and the observed hosts to be primarily due to a difference in stellar age in modeled spectra. The median inferred spectra in Ding et al. (2023) uses a stellar population with an age of 370 Myr and 500 Myr for the galaxies of J2236+0032 and J2255+0255, respectively. For the primary BlueTides analogs, the average age is approximately 100 Myr resulting in a much steeper UV slope. However, the BlueTides galaxy spectra appear broadly consistent with the inferred spectra from the lower age range modeled in Ding et al. (2023).
| J2255 | F356W | 24.2 | 12.5 | ||
| J2255 | F150W | 24.8 | 14.7 | ||
| J2236 | F356W | 23.5 | 18.9 | ||
| J2236 | F150W | 24.6 | 15.0 |
5.2 Mock Rest Frame Images of BlueTides Analogs
With the intrinsic properties simulated in BlueTides, we also create rest frame images of galaxies. These provide a glimpse into the evolution of detected hosts of the Ding et al. (2023) and other SHELLQs quasars that is unachievable with observations. We use SynthObs to generate rest frame images of the BlueTides analogs. These images do not have the observational effects of the other mock images, such as noise and PSF effects, and have a much higher resolution. We show the stellar mass distribution and recent star formation in Figure 6. We also show images in idealized top hat filters in the far UV, visible (V), and infrared (H) as well as the stacked RBG version. These images provide a higher resolution snapshot of the star particles in the BlueTides analog than the blurred redshifted images shown in Figures 7(a) through 8(b). The galaxies are notably compact in the left three panels in stellar mass, SFR, and intrinsic FUV. The intrinsic FUV panel traces the star formation and stellar mass. This is in line with results from Marshall et al. (2022a) which predicted that galaxies hosting quasars are more compact intrinsically relative to the entire BlueTides sample. When we include dust attenuation in the four right panels of Figure 6, we see a flatter emission profile as predicted in Marshall et al. (2022a). This results in the appearance of an extended galaxy. The ring around the center of these galaxies shows a region where the dust is particularly strong.
@@ 5.3 Point source removal
After the creation of mock images described in section 4.3, we compare the mock imaged BlueTides galaxies before and after the point source removal process to further increase the confidence that the Ding et al. (2023) hosts appear to be plausible detections. To do this, we perform quasar subtraction using the software psfmc on our model images, using a mock star image as our input PSF model. For our mock star image, we use the same model PSF from WebbPSF used to create our quasar image but with a different realization of the background and shot noise (see Marshall et al. (2021)). This is an idealized case, where the PSF shape is known perfectly up to noise variation. While the most significant challenge in high-z quasar host detection is often the PSF mismatch error, our approach shows a more ideal case, which is useful for understanding the optimal outcome from these subtractions. Even with a perfect PSF shape, we will inevitably end up with some PSF subtraction error, which we begin to understand qualitatively here, and our aim is to see how this might appear in the Ding et al. (2023) host images.
We recover the properties of the host galaxy and quasar from the imaged galaxies using the software psfmc to do point source extraction. psfmc was written specifically for quasar subtraction unlike other algorithms that perform similar 2D surface brightness fitting such as galfit in Peng et al. (2002). psfmc uses the Affine Invariant Markov chain Monte Carlo (MCMC) ensemble sampler implemented in emcee (Foreman-Mackey et al., 2013). The posterior sampled includes priors specified by the user and the following Gaussian joint likelihood for all pixels
@@
The output of psfmc contains the posteriors of all parameters and the final galaxy image with the quasar subtracted. The intermediate images derived are the raw model and the convolved model. The raw model contains the highest probability galaxy and quasar surface brightness distribution. We convolve the raw model with the PSF to form the convolved model which is subtracted from the input observed image to determine the residuals. The third column of the panels labeled "model" in figures 7(a) through 7(d) show the convolved model.
@@
We initiate 64 walkers and discard the initial 200-400 samples to allow proper convergence of our sampler. The values of our results are drawn from the last 1k sample ensuring accurate parameter error estimation. The mock images are shown and compared qualitatively in the subsequent sections 5.4 and 5.5.
@@ 5.4 Mock Images of BlueTides Analogs
In Figures 7(a) through 8(b), we show six panels of our mock observed images, depicting I) the BlueTides galaxy without a quasar imaged, II) the same BlueTides galaxy with a quasar imaged, III) the convolved model from psfmc described in section 4.3, IV) the host galaxy image with the quasar subtracted via psfmc, V) a normalized residual between the data and point source subtracted image from psfmc (i.e., (data - model) / ), and VI) the detected host JWST image from Ding et al. (2023). With psfmc, we are successful in extracting the 2D galactic brightness distribution and quasar parameters for all of our imaged BlueTides galaxies. This is promising as the analogs are particularly faint in the F150W filter. The actual host (galaxy only image) versus quasar subtracted (data - quasar image) all recover the underlying host with moderate visible error. In Figures 7(b) and 7(c), the center of the host is missing flux. Each quasar subtracted host also features subtle to significant morphological changes from the actual host image. Through the normalized residuals in the fourth panel, we can quantitatively compare the data + quasar versus galaxy only images which ideally would be the same. There is a slight underextraction and overextraction around the core area containing the point source. This is an artifact of the point source removal and should be considered with point source subtraction on real images.
In Figures 7(a) through 7(c), the J2236+0032 and J2255+0251 host analogs are similar to the detected hosts in Ding et al. (2023) in all filters. In Figure 7(d), the J2255+0251 host analog detection in F150W is shown alongside the non-detection in Ding et al. (2023). Notably, this is also the dimmest host analog detection in our sample.
@@ 5.5 Measured Properties in Mock Images
Qualitatively, the imaged galaxies look similar to the detected galaxies in Ding et al. (2023), with similar shapes and compactness. However for most of these mock images, our point source subtraction algorithm described in section 4.3 creates moderate to large errors in predicted radii and magnitude when extracting a point source. We compare the Sérsic radii and magnitudes quantitatively for the imaged galaxy with and without the point source. We extract both Sérsic radii and magnitudes as defined in equation 1 using the MCMC framework in psfmc to maintain consistency. We tabulate the mean and one sigma value from our sampled posteriors for each Sérsic radius in Table 4 for both the original galaxy before adding the point source and the radius estimated after point source subtraction. We find that the results do not always agree to one sigma for both the original and point source subtracted galaxies indicating that point source subtraction is introducing bias.
@@
We note that the errors we find for our analogs are smaller than those seen for the observed quasars and their hosts in Ding et al. (2023). Our errors include errors from our more idealized point source removal, as well as the noise in the image. In the observations in Ding et al. (2023), the uncertainty also includes larger systematic errors due to PSF mismatching and other effects caused by the real telescope imaging process.
@@ 6 Detection predictions for J2255+0251 and J2236+0032
Follow up observations with JWST of J2255+0251 and J2236+0032 (detected in both F150W and F356W by JWST) will allow further constraints on the properties of their hosts. In JWST Cycle 2 GO 3859 (PI: Onoue), these quasars will be followed up in five NIRCam filters. In Table LABEL:tab:jwst_proposal_2, the filters and exposure times for this proposal are shown. In Figure 10, we show the mock images of galaxies without an added quasar for these filters and exposure times. We also extend the exposure time of the original filters (F356W and F150W) to 10ks in Figure 11 for comparison to the 3.1ks exposure times shown in figures 7(a) through 7(d). We supersample in all filters with wavelengths larger than (filters above F200W) to maintain consistency with our mock imaging in section 4.